# RedNote-Vibe: A Dataset for Capturing Temporal Dynamics of AI-Generated Text in Social Media

### v1

@@ Abstract
The proliferation of Large Language Models (LLMs) has led to widespread AI-Generated Text (AIGT) on social media platforms, creating unique challenges where content dynamics are driven by user engagement and evolve over time. However, existing datasets mainly depict static AIGT detection. In this work, we introduce RedNote-Vibe, the first longitudinal (5-years) dataset for social media AIGT analysis. This dataset is sourced from Xiaohongshu platform, containing user engagement metrics (e.g., likes, comments) and timestamps spanning from the pre-LLM period to July 2025, which enables research into the temporal dynamics and user interaction patterns of AIGT. Furthermore, to detect AIGT in the context of social media, we propose PsychoLinguistic AIGT Detection Framework (PLAD), an interpretable approach that leverages psycholinguistic features. Our experiments show that PLAD achieves superior detection performance and provides insights into the signatures distinguishing human and AI-generated content. More importantly, it reveals the complex relationship between these linguistic features and social media engagement. The dataset is available at https://github.com/testuser03158/RedNote-Vibe.
@@ 1 Introduction
Large Language Models (LLMs) (Achiam et al., 2023; Mallick & Kilpatrick, 2025; Guo et al., 2025) have revolutionized digital content creation, leading to a surge in AI-Generated Text (AIGT). Consequently, developing robust detection methods has become a critical research frontier, with a focus on classification (Gui et al., 2025; Hu et al., 2023) and source attribution (Sun et al., 2025). However, existing research primarily treats AIGT detection as a static classification task on formal corpora (e.g., news, academic writing), a paradigm that misaligns with the unique ecosystem of social media.
@@
In this work, we identify two critical and unaddressed challenges for AIGT research on social media. First, unlike formal text, where factual accuracy is prioritized, social media ecosystem rewards content that maximizes user engagement, such as likes, comments, and shares (Chung et al., 2023; Cascio Rizzo et al., 2024). On social media platforms that value sharing real-life experiences, LLMs can be prompted to generate sensationalized or controversial content to increase interactions, which undermines connections and trust within the community. Second, over time, AI and human content are interacting more frequently on social media platforms. This creates a potential co-evolution of linguistic styles, which may also influence community topics, user behavior, and engagement patterns.
Existing datasets and research paradigms typically treat AIGT detection as a static classification task on a fixed snapshot of data, fail to capture these longitudinal trends or explain the relationship between linguistic features and engagement metrics. This research gap is not merely a technical oversight but an approaching cultural issue, as the social media ecosystem could be transforming rapidly. Based on this, measuring and identifying these dynamics is of paramount importance.
@@
To bridge this gap, we introduce RedNote-Vibe, the first dataset designed for studying AIGT in a dynamic social media context. RedNote-Vibe is collected from Xiaohongshu (RedNote), a leading Chinese social media platform. Each sample is enriched with metadata including topic, tags, timestamp, user engagement metrics (i.e., likes, comments, collections) and their parallel AIGT variants generated by a diverse set of LLMs. Notably, our data collection spans a wide timeline, covering content from before the release of ChatGPT to the present (July 2025). This provides a natural testbed for researchers to observe the evolution and impact of LLMs within social media environments.
To address the limitations of existing detection models, which lack the necessary interpretability to link linguistic artifacts to user engagement, we propose the PsychoLinguistic AIGT Detection Framework (PLAD). PLAD first quantifies text into a suite of psycholinguistic features and then utilizes a decision tree-based model for classification. Our experiments show that PLAD not only achieves superior performance with model-based methods but also offers clear insights into the stylistic signatures of different LLMs and reveals how AIGT correlates with user engagement. The main contributions of this paper are as follows:
  • We introduce RedNote-Vibe, a social media AIGT dataset featuring rich engagement metadata and a longitudinal timeline, enabling research on temporal dynamics and user interaction patterns.
  • We propose PLAD, an interpretable, psycholinguistic-based framework that offers strong detection performance while illuminating the connection between linguistic style and social media engagement.
@@
  • We provide comprehensive analysis of our dataset, uncovering temporal trends in AI adoption and revealing differences in engagement patterns between human-authored and AI-generated content.
@@ 2 RedNote Engagement Dataset
RedNote (Xiaohongshu) [^1] https://www.xiaohongshu.com stands as one of the most influential Chinese social media platforms, serving over 300 million monthly active users. This platform emphasizes personal experiences and lifestyle sharing, which makes it particularly vulnerable to AIGC infiltration, as it undermines the authenticity of the content. Despite its influence, RedNote has remained largely unexplored in academic research due to the absence of a publicly available dataset. In this section, we present our RedNote-Vibe dataset, the first large-scale dataset from this platform that captures both temporal dynamics and engagement patterns, specifically designed for research on AIGT detection and the impact of AIGC on social media.
@@ 2.1 Data Collection and Statistics
Our data collection methodology is grounded in RedNote’s official user behavior report, which identifies ten dominant content categories: Career, Wellness, Travel, Health, Food, Pets, Education, Sports, Fashion, and Relationships. We first extract the example tags provided in the report for each category, then expand them to approximately 50 representative tags per topic through manual curation. These expanded tag sets serve as our retrieval queries to ensure comprehensive coverage of each domain.
We adopt a web crawler to collect 120,000 notes from January 2020 to July 2025. To ensure accurate topic classification, we filter these notes using Qwen-2.5-7B, resulting in 98,714 notes. Each sample contains comprehensive metadata including: 1) Content: note title, text content and tags; 2) Temporal information: publication timestamp; 3) Engagement metrics: likes, comments and collections; 4) Topic domain.
Table 1 presents detailed statistics across domains, which shows that different domains exhibit distinct linguistic properties (e.g., average length) and engagement dynamics. Figure 1 visualizes the distribution of the total engagement (defined as the sum of all three metrics), which follows a long-tail pattern consistent with real-world social media. As shown in the inner ring of the chart, a substantial proportion of posts receive very low engagement (0-10 interactions), indicating that most user-generated content attracts limited audience attention. In contrast, a small fraction of posts achieve disproportionately high engagement.
Domain # Length Likes Comm. Colls.
Health 14.9k 398.0 383.3 38.8 322.0
Fashion 11.4k 273.0 718.6 40.9 265.9
Food 7.9k 427.0 49.0 11.2 35.5
Career 11.1k 485.8 381.5 72.6 294.6
Pets 2.1k 457.0 33.9 11.0 18.9
Education 4.8k 529.8 47.0 5.9 33.5
Sports 4.4k 559.6 38.6 9.4 21.3
Relation. 15.5k 372.5 522.2 171.9 183.8
Travel 10.5k 549.0 373.4 59.8 295.4
Wellness 10.7k 483.6 552.5 110.9 282.1
Figure 1: The distribution of posts by domain and interaction count (sum of likes, comments and collections).
@@ 2.2 AI Samples Generation
To create a comprehensive AIGT detection benchmark, we construct parallel AI-generated versions of RedNote posts. Our generation protocol employs a seed-based approach: each human-written note serves as a reference for LLMs to generate stylistically similar but semantically distinct content. Specifically, we provide LLMs with the original note’s title, content, and domain classification, then instruct them to create new posts that: (1) emulate the writing style and tone of the reference; (2) preserve personal characteristics such as colloquialisms, punctuation patterns, and occasional grammatical imperfections typical of social media; (3) maintain comparable text length; and (4) avoid direct copying of phrases or sentences from the original post.
@@
Figure 2 illustrates our data construction pipeline and the selected LLM ensemble. This approach yields a training and validation set with verified human/AI labels. Additionally, we compile an exploration set containing posts from the post-LLM period (2023-2025). While lacking ground-truth labels, this subset enables researchers to investigate real-world content evolution and analyze the emerging linguistic landscape shaped by widespread AI adoption.
@@ 2.3 Task Definition
Figure 2: Overview of the RedNote-Vibe dataset construction. (1) Human-Generated Set is collected in the pre-LLM period before 2022.12, therefore labeled as human-authored. (2) AI-Generated Set is created by prompting a diverse LLM pool with the human seeds to produce synthetic variants. (3) Exploration Set contains post-LLM social posts, enabling AIGT detection and extensive temporal and cross-sectional analyses comparing AI and human content.
@@ 3 PsychoLinguistic AIGT Detection Framework
To address the unique challenges of AIGT detection in dynamic social media environments, we propose the PsychoLinguistic AIGT Detection Framework (PLAD). Unlike existing methods that often fail to explain why a text is classified as AI-generated, PLAD leverages established psychological theories to create an interpretable detection system that not only distinguishes AI from human content but also reveals the underlying linguistic mechanisms.
3.1 PsychoLinguistic Feature Framework
Our framework quantifies a total of 31 linguistic features around four dimensions of human language expression, each rooted in psychological and cognitive theories. The definition of dimensions is shown below:
Emotional and Social Grounding. Human communication is deeply rooted in personal experience and social awareness. Authentic emotional expression is often linked to memory and is conveyed through rich sensory details Conway & Pleydell-Pearce (2000). Furthermore, humans naturally adapt their language to their audience, demonstrating an implicit Theory of Mind through markers of empathy and social connection (Baron-Cohen, 1997). We capture these aspects through features measuring emotional intensity, personal grounding, social connection, and specific markers such as emoji usage patterns (Felbo et al., 2017).
Cognitive Architecture. Drawing from conceptual complexity theory (Baker-Brown et al., 1992) and narrative structure research (Labov & Waletzky, 1997), human authors exhibit multiperspectival reasoning and tolerance for ambiguity. Human argumentation characteristically incorporates nuanced counterarguments (Toulmin, 2003) and embeds coherent value systems (Graham et al., 2009). This dimension encompasses features that quantify perspectival complexity, dialectical reasoning, and temporal coherence.
Lexical Identity and Stylistic Signature. Over time, human writers develop a unique stylistic idiolect (Argamon et al., 2003), characterized by word choices, rhythmic patterns, and natural imperfections such as hesitations, self-corrections, and topic shifts (Clark & Tree, 2002), which signal real-time cognitive processing. While LLMs can imitate various styles, their output exhibits a stochastically uniform distribution of words that reveals their non-human origin. This dimension focuses on quantifying the uniqueness of the lexicon, the stylistic consistency, and the presence of natural linguistic imperfections that signal the human author.
Cohesion and Textual Flow. While the previous dimension assesses word choices, this dimension evaluates the organization and progression of the full content. The human composition naturally evolves through interconnected ideas, exhibiting dynamic semantic progression and adaptive referential chains (Halliday & Hasan, 2014). In contrast, AI-generated text often maintains paragraph-level fluency while lacking deep thematic development, producing semantically static content. We also measure repetition patterns: humans employ strategic repetition for emphasis or clarification (Stamatatos, 2009), which is rarely seen in AI text.
3.2 Feature Extraction and Classification
@@
Our feature set contains 31 features that can be categorized into two extraction approaches. (1) Directly computable statistical features that can be obtained using straightforward computational methods, such as emoji density, type-token ratio and other structural measures. (2) Semantically-based features that require evaluation criteria and text analysis tools for assessment. In contrast to traditional approaches to psychological text analysis such as LIWC (Pennebaker et al., 2015), which employs frequency-based word analysis, we adopt a more sophisticated approach using a proxy LLM for psychological text analysis.
Specifically, following related work (Rathje et al., 2024; Ghatora et al., 2024), we design evaluation rubrics that convert theoretical constructs into measurable criteria for characteristics. These rubrics are then presented to the proxy LLM, who is instructed to evaluate the input text according to the specified dimensions and provide quantitative scores. To mitigate potential biases introduced by proxy models, all results undergo a verification mechanism using Chain-of-Thought reasoning to ensure accuracy. Experimental validation demonstrates that our configuration achieves higher correlation with human annotations compared to existing methods. The detailed feature list and example of criteria are shown in the Appendix.

### v2

# RedNote-Vibe: A Dataset for Capturing Temporal Dynamics of AI-Generated Text in Lifestyle Social Media

@@ Abstract
We introduce RedNote-Vibe, a dataset spanning five years (pre-LLM to July 2025) sourced from lifestyle platform RedNote (Xiaohongshu), capturing the temporal dynamics of content creation and is enriched with comprehensive engagement metrics. To address the detection challenge posed by RedNote-Vibe, we propose the PsychoLinguistic AIGT Detection Framework (PLAD). Grounded in cognitive psychology, PLAD leverages deep psychological signatures for robust and interpretable detection. Our experiments demonstrate PLAD’s superior performance and reveal insights into content dynamics: (1) human content continues to outperform AI in emotionally resonant domains; (2) AI content is more homogeneous and rarely produces breaking posts, however, this human-AI gap narrows for arousing higher-investment interactions; and (3) most interestingly, a small group of users who strategically utilize AI tools can achieve higher engagement outcomes. https://github.com/ydli-ai/RedNote-Vibe
RedNote-Vibe: A Dataset for Capturing Temporal Dynamics of AI-Generated Text in Lifestyle Social Media
Yudong Li1   Yufei Sun 2   Peiru Yang1*   Yuhan Yao2   Wanyue Li3   Jiajun Zou1 Haoyang Yang4Haotian Gan1Linlin Shen4Yongfeng Huang1 1Tsinghua University  2Beijing University of Posts and Telecommunications 3 Hong Kong Metropolitan University  4Shenzhen University liyudong@tsinghua.edu.cn, ypr21@mails.tsinghua.edu.cn
@@ 1 Introduction
Large Language Models (LLMs) (Achiam et al., 2023; Mallick and Kilpatrick, 2025; Guo et al., 2025) have revolutionized digital content creation, leading to a surge in AI-Generated Text (AIGT). Consequently, developing robust detection methods has become a critical research frontier, with a focus on classification (Gui et al., 2025; Hu et al., 2023) and source attribution (Sun et al., 2025a). However, existing research primarily focus on knowledge-sharing and formal corpora (e.g., news, academic writing). There remains a lack of research on lifestyle-sharing social media platforms, where content is deeply intertwined with personal narratives and commercial influence.
@@
In this work, we identify two critical and unaddressed challenges for AIGT research on social media. First, unlike formal text, where factual accuracy is prioritized, social media rewards content that maximizes user engagement, such as likes, comments, and shares (Chung et al., 2023; Cascio Rizzo et al., 2024). On social media platforms that value sharing real-life experiences, LLMs can be prompted to generate sensationalized or controversial content to increase interactions, which undermines connections and trust within the community. Second, over time, AI and human content are interacting more frequently on social media platforms. This longitudinal mixture of content sources presents challenges for detection methods, as linguistic patterns may shift over time. Existing datasets, typically static snapshots from informational sources (e.g., Reddit, Quora), fail to capture these temporal dynamics or the engagement patterns of commercialized social content.
@@
To bridge this gap, we introduce RedNote-Vibe, the first dataset designed for studying AIGT in a dynamic, lifestyle-centric social media context. RedNote-Vibe is collected from RedNote (Xiaohongshu), a leading Chinese platform combining social networking with e-commerce. Our data spans a wide timeline (Oct 2020 - July 2025), covering the transition from the pre-LLM era to the post-LLM era. The dataset features: (1) rich metadata including engagement metrics (likes, comments, collections); (2) parallel AIGT variants generated by 17 diverse LLMs based on human seeds; and (3) an exploration set of real-world post-LLM content. It provides a natural testbed for observing the evolution of AIGT and its impact on user interaction.
Furthermore, facing the challenge that as LLMs evolve, surface-level linguistic artifacts (e.g., syntax patterns) will inevitably be neutralized. We propose the PsychoLinguistic AIGT Detection Framework (PLAD). Instead of relying solely on statistical patterns, PLAD leverages linguistic features grounded in psycholinguistic theories to construct interpretable detection models.
Our experiments demonstrate PLAD not only achieves superior performance compared to black-box baselines but also offers insights into how specific psychological signatures correlate with social media engagement. The main contributions of this paper are as follows:
Figure 1: RedNote-Vibe dataset distribution.
  • We introduce RedNote-Vibe, a large-scale Chinese social media AIGT dataset. To our knowledge, it is the first dataset from the RedNote platform.
  • We propose PLAD, an interpretable framework leveraging psycholinguistic theories, offering robust detection and explaining the linguistic style of AIGT.
@@
  • We provide a comprehensive analysis of the dataset, uncovering temporal trends in AI adoption and revealing how AI-generated content differs from human authorship in driving social interactions.
@@ 2 RedNote Engagement Dataset
Figure 2: Overview of the RedNote-Vibe dataset construction. (1) Human-Generated Set is collected in the pre-LLM period before 2022.12, therefore labeled as human-authored. (2) AI-Generated Set is created by prompting a diverse LLM pool with the human seeds to produce synthetic variants. (3) Exploration Set contains post-LLM social posts, enabling AIGT detection and extensive analyses comparing AI and human content.
RedNote (Xiaohongshu) [^1] https://www.xiaohongshu.com stands as one of the most influential Chinese social media platforms, serving over 300 million monthly active users. This platform emphasizes personal experiences and lifestyle sharing, which makes it particularly vulnerable to AIGC infiltration, as it undermines the authenticity of the content. Despite its influence, RedNote has remained largely unexplored in academic research due to the absence of a publicly available dataset. In this section, we present our RedNote-Vibe dataset, the first publicly available dataset from this platform. It also captures temporal dynamics and engagement patterns, specifically designed for research on AIGT detection and the impact of AIGC on social media.
Domain # Length #Tags Likes Comm. Colls.
Health 19.2k 333.6 7.6 2814.3 97.1 2468.8
Fashion 17.7k 191.2 10.9 1850.5 62.5 681.3
Food 10.5k 398.4 3.0 66.2 11.9 45.0
Career 11.7k 406.9 10.0 2047.2 102.6 1445.9
Pets 3.9k 433.2 2.7 57.5 12.6 34.1
Education 5.9k 496.0 3.7 55.9 6.1 38.3
Sports 17.7k 528.8 3.2 60.1 9.8 36.4
Relation. 22.5k 308.2 7.1 2663.6 275.8 593.9
Travel 14.8k 452.8 10.7 1102.4 86.6 662.0
Wellness 16.0k 396.5 10.5 2348.6 217.7 1025.8
Others 16.9k 220.9 8.9 208.8 1266.2 3225.0
@@ 2.1 Data Collection and Statistics
Our data collection methodology is grounded in the 2024 RedNote User Behavior Report  [^2] http://xhslink.com/o/3ZtMUSbgoCl, which identifies ten dominant content categories: Career, Wellness, Travel, Health, Food, Pets, Education, Sports, Fashion, and Relationships. Our dataset focuses on the Chinese language. To ensure our dataset reflects real-world distribution, we employ a two-step retrieval strategy: We first extracted the official popular search keywords provided in the report for each category. For instance, the Relationships category includes specific seed tags such as “Should couples go dutch?” (情侣之间应该AA制吗) and “First meeting with a blind date” (相亲第一次见面). Then, we expand these tag lists based on search recommendations, resulting in about 500 tags for different domains. We use these tags as search keywords to collect the notes.
We collected 148,000 notes published between October 2020 and July 2025. To maintain the quality of the text content, we filter out notes with fewer than 100 Chinese characters, which were mainly image-based. Finally, our data source contains samples with comprehensive metadata including: 1) Content: note title, text content and tags; 2) Temporal information: publication timestamp; 3) Engagement metrics: likes, comments and collections; 4) Topic domain.
Table 1 presents detailed statistics across domains, which shows that different domains exhibit distinct properties and engagement dynamics. Fig. 1-(a) visualizes the distribution of the total engagement (defined as the sum of all three metrics), Fig. 1-(b) presents the distribution of time span. It can be seen that our dataset covers different levels of interaction metrics, reflecting real-world user interaction behavior on social media.
Dataset Data Source(s) # Time Span Timestamps Language(s) LLMs Interact.
SM-D (2025b) Medium, Quora, Reddit 2.4M Jan 2022 - Oct 2024 English 12
RAID (2024) News, Books, Reddit, etc. 6.2M Not Applicable English 11
MultiSocial (2024) Twitter, WhatsApp, etc. 472k Mostly Pre-2022 Multi (22) 7
MULTITuDE (2023) News 74k Pre-2021 Multi (11) 8
SAID (2023) Zhihu 72k Mar 2023 - Oct 2023 Chinese
RedNote-Vibe RedNote 148k Oct 2020 - Jul 2025 Chinese 18
Table 2: Comparison of RedNote-Vibe with prior social media AIGT datasets. Interact.: SM-D provides the number of followers, comments and likes.
@@ 2.2 AI Samples Generation
To create a comprehensive AIGT detection benchmark, we construct parallel AI-generated versions of RedNote posts. Unlike existing AIGT datasets that mainly direct prompt LLMs to generate AI samples, we employ a pipeline approach to simulate AIGT challenges in real-world scenarios. We design the generation method to be difficult for humans to distinguish. The detailed prompt is shown in Appendix C.1.
@@
Figure 2 illustrates our data construction pipeline and the selected LLM pool. This approach yields a training and validation set with verified human/AI labels. Additionally, we compile an exploration set containing posts from the post-LLM period (2023-2025). While lacking ground-truth labels, this subset enables researchers to investigate real-world content evolution and analyze the emerging linguistic landscape shaped by widespread AI adoption.
In Table 2, we compare our dataset with existing publicly available datasets for AIGT detection in social media. While the existing datasets are valuable, RedNote-Vibe offers unique time spans and timestamps that support AIGT evolution and trend studies in Chinese context. Compared with most related SM-D Sun et al. (2025b), its sources are rooted in knowledge sharing and community discussions (e.g., Quora, Reddit), RedNote is a lifestyle-sharing platform integrated with commercial activities. This presents distinct challenges: AIGT on RedNote is often crafted into personal narratives (e.g., product reviews, travel experiences, etc.), making it significantly harder to detect compared to informational text. Furthermore, given the platform’s commercial attributes, such content poses a more direct influence on consumer behavior. Therefore, RedNote-Vibe not only fills the gap of data source from the RedNote platform but also provides a novel testbed for studying AI-generated content in high-stakes, commercially driven social environments.
@@ 3 PsychoLinguistic AIGT Detection Framework
Dimension Features & Theoretical Basis
Emotional & Social Grounding
Evaluates emotional depth, social connection, and sensory richness. rooted in personal experience and social awareness.
$\bullet$ Emotional Intensity (Gross, 1998); Personal Emotional Grounding (Conway and Pleydell-Pearce, 2000)
$\bullet$ Sensory Detail Richness (Levine et al., 2002); Social Connectedness (Giles and Powesland, 1975)
$\bullet$ Empathetic Engagement (Baron-Cohen, 1997); Interactive and Dialogic Stance (Bakhtin, 2010)
$\bullet$ Unique Emoji Ratio (Felbo et al., 2017); Emoji Density (Felbo et al., 2017)
Cognitive Architecture
Assesses perspectival complexity, reasoning, and narrative structure.
$\bullet$ Perspectival Complexity (Baker-Brown et al., 1992); Narrative Structure Flexibility (Labov and Waletzky, 1997)
$\bullet$ Dialectical Argumentation Strength (Toulmin, 2003); Self-Correction (Hyland, 2005)
$\bullet$ Axiological Coherence (Graham et al., 2009); Temporal Orientation and Integration (Reichenbach, 1947)
$\bullet$ Sentence/Word/Character Count (Manning and Schutze, 1999)
Lexical Identity & Stylistic Signature
Captures stylistic idiolect, rhythmic patterns, and natural imperfections.
$\bullet$ Lexical-Stylistic Personalization (Argamon et al., 2003); Prosodic Rhythm Consistency (Halliday and Matthiessen, 2013)
$\bullet$ Imperfection (Clark and Tree, 2002); Rhetorical Sophistication (Grice, 1975)
$\bullet$ Punctuation Ratio, Number Ratio (Manning and Schutze, 1999)
$\bullet$ Type-Token Ratio (Templin, 1957); Word Frequency Entropy (Shannon, 1948); Word Burstiness (Gries, 2008)
Cohesion & Textual Flow
Evaluates the organization, semantic progression, and repetition patterns.
$\bullet$ Lexical Cohesion (Halliday and Hasan, 2014); Inter-Sentential Sentence Similarity (Foltz et al., 1998)
$\bullet$ Immediate Repetition Density, Phrasal Repetition Frequency (Stamatatos, 2009)
$\bullet$ Sentence Burstiness (Gries, 2008)
Table 3: Feature list of the PLAD framework, categorized by psychological dimensions. Each feature is grounded in established psycholinguistic theories.
Existing approaches to AIGT detection primarily rely on computational linguistic features or supervised training of PLMs, achieving significant performance. However, as AI continues to evolve, the linguistic features such as style and syntax that currently distinguish AIGT will inevitably be neutralized. We hypothesize that detection methods grounded in established psycholinguistic dimensions (e.g., emotional expression patterns, narrative coherence, and lexical diversity) may be more robust to model updates, as these dimensions reflect fundamental differences in how text is generated. Consequently, we propose the PsychoLinguistic AIGT Detection Framework (PLAD), which leverages constructs from psycholinguistics as measurable features for interpretable and potentially more stable AIGT detection.
Our framework quantifies a total of 31 features categorized into four dimensions of human language expression: Emotional and Social Grounding, Cognitive Architecture, Lexical Identity and Stylistic Signature, and Cohesion and Textual Flow. These dimensions are rooted in psychological and cognitive theories, capturing aspects from emotional depth to narrative complexity. Table 3 presents the definition of each dimension and the corresponding features with their references.
@@
Our feature set can be categorized into two extraction approaches. (1) Directly computable statistical features that can be obtained using straightforward computational methods, such as emoji density, type-token ratio and other structural measures. (2) Semantically-based features that require evaluation criteria and text analysis tools for assessment. In contrast to traditional approaches to psychological text analysis such as LIWC (Pennebaker et al., 2015), which employs frequency-based word analysis, we adopt a more sophisticated approach using a proxy LLM for psychological text analysis. Specifically, following related work (Rathje et al., 2024; Ghatora et al., 2024), we design evaluation rubrics that convert theoretical constructs into measurable criteria for characteristics. These rubrics are then presented to a proxy LLM, which is instructed to evaluate the input text according to the specified dimensions and provide quantitative scores.
Considering that using a proxy LLM for feature extraction may bring model biases into the feature set. However, we adopt this approach based on several reasons: (1) For complex psycholinguistic constructs, LLMs currently represent the most versatile and scalable assessment method; (2) Recent studies have demonstrated the robustness of LLM-based psychological assessment, and our validation of LLMs against human annotations aligns with previous results. The detailed implementation and evaluation are shown in Appendix C.